# Mastering Board Games by External and Internal Planning with Language Models

### v1

@@ Abstract
While large language models perform well on a range of complex tasks (e.g., text generation, question answering, summarization), robust multi-step planning and reasoning remains a considerable challenge for them. In this paper we show that search-based planning can significantly improve LLMs’ playing strength across several board games (Chess, Fischer Random / Chess960, Connect Four, and Hex). We introduce, compare and contrast two major approaches: In external search, the model guides Monte Carlo Tree Search (MCTS) rollouts and evaluations without calls to an external engine, and in internal search, the model directly generates in-context a linearized tree of potential futures and a resulting final choice. Both build on a language model pre-trained on relevant domain knowledge, capturing the transition and value functions across these games. We find that our pre-training method minimizes hallucinations, as our model is highly accurate regarding state prediction and legal moves. Additionally, both internal and external search indeed improve win-rates against state-of-the-art bots, even reaching Grandmaster-level performance in chess while operating on a similar move count search budget per decision as human Grandmasters. The way we combine search with domain knowledge is not specific to board games, suggesting direct extensions into more general language model inference and training techniques.
@@ 1 Introduction
With the recent advances in large language models (LLMs) and their widespread adoption across diverse application domains and use cases, their capabilities are under scrutiny. While LLMs perform fluently on text generation, language understanding and translation, they are prone to hallucinations and reasoning errors, especially in complex contexts (Chang et al., 2024; Hadi et al., 2023). Hence, special attention has been given to the development of planning and reasoning capabilities in LLMs (Minaee et al., 2024), and/or to the agentic systems built on top of LLM subcomponents (Wang et al., 2024e; Grosnit et al., 2024). This reflects both the increasing complexity of AI application domains, as well as a recognition of reasoning as a key capability on the path towards more general artificial intelligence (Huang and Chang, 2022). In terms of Kahneman’s cognitive theory of two systems (Kahneman, 2011), prior work primarily improved associative System 1 inference in language models, whereas planning and reasoning now focuses on improving the more deliberate System 2 thinking (Plaat et al., 2024).
Reasoning in language models encompasses diverse techniques with the aim to improve LLM performance on reasoning benchmarks. Approaches typically fall into one of two distinct categories: In internal planning the LLM develops a plan in context, like Chain-of-Thought prompting (Wei et al., 2022), by autoregressively considering possible steps towards the goal and their consequences. By contrast, external planning uses the LLM to generate steps in a neurosymbolic system, such as in Tree of Thought (Yao et al., 2024), where an outer loop performs explicit search over possible sequences of steps. This paper presents how language models can be trained for internal and external planning to improve reasoning in sequential decision-making, using board games as an experimental domain.
@@
First, we have pre-trained a Transformer model, the multi action-value (MAV) model, capable of playing several board games (Chess, Fischer Random / Chess960, Connect Four, Hex) at a strong level. Our evaluation shows that this model is capable of reliably tracking the board state throughout games, and makes (good) legal moves. We observe nearly no hallucinations (cf. Table 3) with this model for both in and out-of-distribution positions that are fairly distinct from anything the model would have seen during training. Our MAV model builds on top of prior work (Ruoss et al., 2024) where the ability to train strong text-based Transformer models in chess without search was first demonstrated, though those methods would still require an external tool for providing the list of legal moves, querying for their values individually. Our MAV model generates all the values and transitions in a single query, and proposes the list of candidate moves itself without any external tool calls. Second, we demonstrate that it is possible to utilize the proposed MAV model within an external MCTS controller, as a value/implicit policy and transition function. The MCTS-MAV agent shows substantial improvements over the base MAV model in playing strength, showcasing the promise of MCTS as an external search mechanism in a more general context. Our asynchronous MCTS implementation has been designed to reduce the number of queries to the model, and we show a comprehensive evaluation of different variants and scoring methods, in the low-simulation regime ($\sim$100 to $\sim$1k simulations compared to AlphaZero which used $\sim$10k simulations or engines that use up to millions). Third, we generalise the approach recently proposed in Stream of Search (SoS) (Gandhi et al., 2024b) and show that it is possible to distill the search procedure directly into the LLM. The resulting internal search MAV model has the ability to infer the distilled search procedure in chess. Our evaluation shows that this method is also capable of delivering consistent improvements in playing strength over the MAV baseline with the given search/token budget. By training on tokens representing various search actions (such as expand, evaluate, summarize) and game actions (moving the state ahead), we show a significant improvement when comparing against training on game actions only, elevating the model’s skill from imitating search to executing it. We summarize below our main contributions:
Contribution 1: MAV model. We pre-train a Transformer model, the multi action-value model (MAV), capable of playing several board games (Chess, Chess960, Connect Four, Hex) at a strong level.
Contribution 2: External search. We use the MAV model within an external MCTS controller, as a value/implicit policy and transition function. The agent reaches Grandmaster level with the number of moves considered per decision comparable to human players.
Contribution 3: Internal search. We distill directly the search procedure into the LLM. The resulting agent is able to infer the search procedure and its performance scales with the given search budget.
@@ 2 Multi-action-value model
The first contribution of this work is to introduce the multi-action-value (MAV) model, a Transformer model pre-trained exclusively on textual game data that functions simultaneously as a: ($i$) world model, ($ii$) value function, and ($iii$) policy function for multiple perfect-information board games. Acting as world model further requires the following capabilities: ($i$) state-tracking (determining the new state of the game after a move has been played in the previous state), ($ii$) legal move prediction, ($iii$) terminal state detection. To achieve this, the MAV model is trained on examples following a flexible format which provides a strong learning signal and allows for a quality-compute tradeoff at inference time while remaining in-distribution. An example illustrating the MAV format can be found in Figure 1.
@@
The MAV model input starts with a configuration header containing a series of commands that inform how to interpret the inputs that follow, and what outputs are expected. In its simplest form, this looks like <mav game=\$game> %state %top_1 </mav>, where %state and %top_1 are commands. This tells the model what game is being played, that the next line ([%state \$state_string]) should be interpreted as the current state of the game, and that it is responsible for predicting the output top_1 (i.e. the single best move and its win probability).
@@
Each game uses a different textual format to represent its state. These representations are designed such that each square of the board is tokenized separately, and squares maintain constant relative positions in token space. This makes it easier for the model to navigate and manipulate the board. This is the format used for the %state and %prev_state commands. For convenience we also support the %FEN and %prev_FEN commands for chess, which instead use the standardized FEN representation. We train our models to be able to handle %state and %FEN commands interchangeably, or jointly as in Figure 1. For simplicity, in the rest of this section we will only use the %state command.
State tracking.
State tracking is achieved in the MAV model by using the series of commands %prev_state %prev_action %state. These instruct the model that the following input lines will be [%prev_state \$state_string] containing the previous state, and [%prev_action \$action_string] with the move played in that state. The model’s task is to act as a transition function and output [%state \$new_state_string], describing the state after the transition.
In addition to predicting the state after the previous action, MAV can also output the state after the action it decides to take (see %best_action below).
@@
When state tracking is used, the post-transition state is used as the starting state for subsequent commands (e.g. %top_1 is evaluated in this state). Otherwise the starting state is directly specified using the %state command.
@@
The %top_k command instructs the model to output a list of $k$ legal moves (or all legal moves if there are fewer than $k$). The ability to configure $k$ is useful because it allows for varying the amount of inference-time computation the MAV model performs. However, it is also possible to specify $k=``all"$ in order for the model to produce all legal moves.
Directly following each legal move the model outputs its action value. This corresponds to the predicted win probability for the player in the current state if this move is taken. For chess, state-action values are represented similarly to Ruoss et al. (2024): Stockfish centipawn evaluations are mapped to win probabilities using the formula
@@
obtained from https://lichess.org/page/accuracy. Win probabilities are then mapped to discrete non-overlapping buckets, making the value prediction problem a classification task rather than a regression task. This has proven to be beneficial in other works (Farebrother et al., 2024). We use 64 buckets, each represented by a different special token (e.g. <ctrl28> for bucket 28). Figure 2 shows two win probability distributions predicted by the MAV model. Connect Four and Hex use game engines Fhourstones (Tromp, ) and neurobenzene (Gao and Pawlewicz, ), respectively, to obtain state-action values, which are similarly mapped to the same 64 bucket tokens. When producing training examples, we randomly vary $k$ and include both examples where $k$ is less than the number of legal moves, and where $k$ is greater (this teaches the model to not hallucinate additional options). In the former case, we teach the model to output the $k$ moves that have the highest win probabilities as predicted by the game engine. We make the deliberate choice to randomize the order of the moves. This accomplishes two purposes: ($i$) it encourages the model to treat the moves independently, rather than relying on a specific order, and ($ii$) it helps to prevent hallucinations. If the moves were e.g. sorted lexicographically from $a1a2$ to $h8h7$, then ($i$) if the model skips a legal move $a1a4$ and emits $a1a5$, there is no hope that it will correct its mistake, and ($ii$) the model might learn to systematically lean on evaluations of previous moves when considering $h8h7$, potentially causing the model to play better/worse on the left side of the board. Using randomised ordering of moves in the training data, the model is steered towards an approximate permutation symmetry (Bronstein et al., 2021).
Figure 2: For a safe positional move, such as Re1 MAV assigns little uncertainty and gives a win probability of 40-45% since white is a pawn down. In contrast, the bishop sacrifice Bxh7+ (best move according to Stockfish) is a riskier attacking move and thus has substantial probability mass around low, equal, and high win probabilities.
@@
As mentioned above, there are 64 buckets used to encode the range of a value, each represented by a unique token. As we reserve these 64 tokens exclusively for encoding values, the actual state-action value for a move is derived from its token distribution. Hence, it can also be viewed as a decoding strategy for the values. We refer to different ways to represent values as scoring methods. We experiment with two scoring methods: max scoring, which corresponds to greedy decoding, uses the mode of model’s distribution over the 64 buckets to represent the value, and mean scoring, which takes into account the whole distribution defining the value as an expectation over buckets.
Mean scoring often enables us to differentiate between multiple candidate actions with equal modes. For example, there are many positions where multiple moves will likely result in a draw but certain moves retain higher winning chances than others. In such cases, max scoring would not be able to differentiate between these moves, while mean scoring would pick the move with the highest expected winning chances.
The mean scoring method was also used in Ruoss et al. (2024); the main difference here is that MAV does not take the move as an input to be scored, but rather decodes the values for all moves in a single output generation.
Legal moves and terminal states.
It’s worth noting that %top_k also handles functionality specifically related to MAV being a world model. First, the model does not need to be provided with a list of legal moves, and instead predicts the list by itself. Second, if the state is terminal, the %top_k output specifies the game outcome (e.g. [%top_1 invalid : "1-0"] when the first player has won).
@@
In positions where the acting player has a decisive advantage, multiple moves may be assigned a 100% win probability, and therefore all map to the same bucket. Some of these moves may not directly lead to finishing the game by checkmate, causing the model to play aimlessly in winning positions. This problem was encountered in Ruoss et al. (2024), which they resolved by using Stockfish to break ties when all top 5 moves lie above a win probability of 99%. We seek to completely eliminate reliance on game engines at inference time. To accomplish this, we introduce the %best_action command, which must follow the %top_k command. If present, the model outputs [%best_action \$move], where \$move corresponds to Stockfish’s chosen action in the position. Essentially, the model learns the tie-breaking procedure described above.
@@
To summarize how the MAV model differs from previous Transformer-based chess engines, the MAV model makes four key improvements over the existing state-of-the-art (Ruoss et al., 2024; Monroe and Leela Chess Zero Team, 2024; Farebrother et al., 2024)
  • First, it performs world modeling, policy and action-value computation together in one model.
  • Second, it is trained to output the best action at the end of action-value modeling. This enables the model to finish games even in the presence of multiple actions with equal values but with some of the actions not bringing the game closer to a finish. Importantly, these first two improvements enable the MAV model to play complete games without relying on an external game engine for legal moves or for finishing the game.
  • Third, all of the above steps can be done in a single model call without having to evaluate every action separately—an important feature that guided the design of the MAV format to reduce the cost and infrastructural complexity of inference.
  • Fourth, the amount of compute performed by the model at inference time can be varied. This enables us to achieve higher quality at the cost of higher latency when performing both internal and external search, where the performance increases as we scale the planning.
@@
We curate a dataset of diverse, relevant positions in four games: Chess, Fisher-Random Chess, Connect Four, and Hex. The statistics and sources for each of these datasets are shown in Table 1. Each position is used to produce a single training example, randomly varying ($i$) the $k$ action values in %top_k, ($ii$) the presence of the initial or final state tracking commands, ($iii$) the use and order of %state or %FEN representations in chess. Further details on the datasets are provided in the Appendix.
Game Positions Action values
Chess 3.1B 54.3B
Chess960 1.2B 20.9B
Connect Four 21.8M 110.5M
Hex 125.6M 537.0M
Table 1: Multi-action-value (MAV) model training data statistics.
@@
We train two randomly initialized decoder-only Transformer models using the Gemini architecture (Gemini Team et al., 2024), called MAV and MAV–small, with 2.7 billion and 1 billion parameters respectively [^1] Note that in practice MAV and MAV–small don’t utilize many natural-language specific parameters.. They were trained on 1.9 and 1.2 epochs of the dataset described in Table 1. Except for the max scoring result for MAV–small in table 2, all other experiments and results use the 2.7 billion MAVmodel. In order to efficiently use our models’ parameters, the input part of each training example is masked out during loss computation. This means our models do not waste capacity on learning to generate game positions.
@@ 3 External Search
External search employs the MAV model to generate planning steps, and applies a search algorithm to direct and optimise over sequences of planning steps. In this paper, we evaluate planning on top of the previously discussed MAV model, using MCTS as the search procedure.
@@
External search is based on an AlphaZero-style MCTS (Silver et al., 2017). There are two key ingredients: a prior function which returns a probability distribution over actions at each state, and a value function returning a numerical value indicating value of a state or state-action pair, both extracted from the MAV output. External search is an adaptation of MCTS; in its most basic form, as in AlphaZero, MCTS relies on an explicit world model, querying a game engine (e.g., a chess implementation in OpenSpiel (Lanctot et al., 2019)) to provide legal actions, state transitions and terminal state signals. Inspired by MuZero (Schrittwieser et al., 2020), in a subsection below we describe how we remove this dependency on the game engine and instead use only the model to track states, transitions, and provide legal actions and game outcome during planning.
The external search algorithm guided by a learned world model is summarized in Algorithm 1, with subroutines and details contained in the Appendix. A search is started at an initial game state $s_{0}$. First, the legal actions $\boldsymbol{a}_{0}^{L}$ and associated state-action values $\boldsymbol{Q}^{(i)}(s_{0},\boldsymbol{a}_{0}^{L})$ from active player’s $i$ perspective are obtained from the MAV model. The prior function is an $\varepsilon$-greedy policy, derived entirely from state-action values comprised of a greedy softmax over the top $k$ values, mixed with a uniform distribution over all actions to encourage exploration. Specifically, let $\boldsymbol{a}^{g,k}\subseteq\boldsymbol{a}^{L}$ be the subset of legal actions $\boldsymbol{a}^{L}$ whose values are among the top $k$-ranked values, or equal to $\boldsymbol{a}^{L}$ if $|\boldsymbol{a}^{L}|\leq k$. Define the probability of action $a$ in state $s$ under the greedy policy to be
$\pi^{g,k}(s,a;\boldsymbol{Q}^{(i)})=\left\{\begin{array}[]{ll}\frac{\exp{\frac{1}{\tau}Q^{(i)}(s,a)}}{\sum_{a^{\prime}\in\boldsymbol{a}^{g,k}}\exp{\frac{1}{\tau}Q^{(i)}(s,a^{\prime})}}&\mbox{if $a\in\boldsymbol{a}^{g,k}$};\\ 0&\mbox{otherwise}.\end{array}\right.$
@@
The probability of the uniform policy is $\pi^{u}(s,a;\boldsymbol{a}^{L})=\frac{1}{|\boldsymbol{a}^{L}|}$. The prior probability of taking action $a$ in state $s$ is then
$\begin{split}P(s,a;\boldsymbol{Q}^{(i)},\boldsymbol{a}^{L},k)&=(1-\varepsilon)\pi^{g,k}(s,a;\boldsymbol{Q}^{(i)})\\ &+\varepsilon\pi^{u}(s,a;\boldsymbol{a}^{L}).\end{split}$ (1)
The root node $\mathcal{N}_{0}$ is then initialized with the active player, string description of a state corresponding to the node, and legal actions (respectively: $\mathcal{N}_{0}(i)$, $\mathcal{N}_{0}(s)$, $\mathcal{N}_{0}(\boldsymbol{a}^{L})$). The node is then expanded so that a child node is added for each action and a prior probability attached to the action leading to the child. Then simulations are run (Algorithm 3 in the Appendix): each simulation starts at the root node, actions and children nodes are selected according to PUCT until a leaf node is expanding and evaluated, and values are backpropagated through the nodes that the simulation visited. After $K$ such simulations, a final move $a^{*}$ is selected. Besides the integration of the prior, value, and transition functions through the MAV model, and the asynchronous simulations described below, this most basic form of external search is based on a fairly common implementation of AlphaZero-style MCTS using PUCT for action selection (Silver et al., 2017) adapted from the generic one in OpenSpiel (Lanctot et al., 2019).

### v3

@@ Abstract
Advancing planning and reasoning capabilities of Large Language Models (LLMs) is one of the key prerequisites towards unlocking their potential for performing reliably in complex and impactful domains. In this paper, we aim to demonstrate this across board games (Chess, Fischer Random / Chess960, Connect Four, and Hex), and we show that search-based planning can yield significant improvements in LLM game-playing strength. We introduce, compare and contrast two major approaches: In external search, the model guides Monte Carlo Tree Search (MCTS) rollouts and evaluations without calls to an external game engine, and in internal search, the model is trained to generate in-context a linearized tree of search and a resulting final choice. Both build on a language model pre-trained on relevant domain knowledge, reliably capturing the transition and value functions in the respective environments, with minimal hallucinations. We evaluate our LLM search implementations against game-specific state-of-the-art engines, showcasing substantial improvements in strength over the base model, and reaching Grandmaster-level performance in chess while operating closer to the human search budget. Our proposed approach, combining search with domain knowledge, is not specific to board games, hinting at more general future applications.
@@ 1 Introduction
While large language models (LLMs) perform fluently on text generation, language understanding and translation, they are prone to hallucinations and reasoning errors, especially in complex contexts (Chang et al., 2024; Hadi et al., 2023). Hence, special attention has been given to the development of planning and reasoning capabilities in LLMs (Minaee et al., 2024). In terms of Kahneman’s cognitive theory of two systems (Kahneman, 2011), prior work primarily improved associative System 1 inference in language models, whereas planning and reasoning now focuses on improving the more deliberate System 2 thinking (Plaat et al., 2024).
Planning and reasoning approaches typically fall into one of two distinct categories: In internal planning the LLM develops a plan in context, like Chain-of-Thought prompting (Wei et al., 2022), by autoregressively considering possible steps towards the goal and their consequences. By contrast, external planning uses the LLM to generate steps in a neurosymbolic system, such as in Tree of Thought (Yao et al., 2024), where an outer loop performs explicit search over possible sequences of steps. This paper presents how language models can be trained for internal and external planning to improve reasoning in sequential decision-making, using board games as an experimental domain.
@@
Contribution 1: MAV model. We pre-train a Transformer model, the multi action-value model (MAV), capable of playing several board games (Chess, Chess960, Connect Four, Hex) at a strong level. This model is capable of reliably tracking the board state throughout games, and makes (good) legal moves.
Contribution 2: External search. We use MAV within an external MCTS controller, as a value/implicit policy and transition function. Our Async MCTS reaches Grandmaster level with the number of moves considered per decision ($\sim$100 to $\sim$1k) comparable to human players (for reference, AlphaZero (Silver et al., 2017) used $\sim$10k simulations and traditional engines use up to millions).
Contribution 3: Internal search. We distill the search procedure directly into the LLM, generalizing the Stream of Search (SoS) (Gandhi et al., 2024b) approach to a significantly more complex domain. The performance of the resulting agent scales smoothly with the given search budget.
@@ 2 Multi-Action-Value Model
The multi-action-value (MAV) model is a Transformer model pre-trained exclusively on textual game data that functions simultaneously as a: ($i$) world model, ($ii$) value function, and ($iii$) policy function for multiple perfect-information board games. Acting as world model further requires the following capabilities: ($i$) state-tracking (determining the new state of the game after a move has been played in the previous state), ($ii$) legal move prediction, ($iii$) terminal state detection. To achieve this, MAV is trained on examples following a flexible format illustrated in Figure 1.
@@
MAV input starts with a header containing a series of commands specifying the game being played, the inputs provided and the expected outputs.
@@
Each game uses a different textual format to represent its state, used with the %state and %prev_state commands. All of them are designed to be easy for the model to navigate and manipulate: each field on the board is a separate token, and fields maintain constant relative positions in token space. For chess we additionally support using the standard FEN representation (with the %fen command).
@@
The starting state for the model can be given in two ways: either directly with the %state command, or as in Figure 1 with the commands %prev_state %prev_action %state, which instruct the model to take a previous state (%prev_state) and a move played in that state (%prev_action) and use these to infer and output the starting state (%state). The latter functionality makes the model act as a transition function.
@@
The %top_k command instructs the model to consider the starting state and output the best $k$ legal moves and their action values, in order of preference from highest action value to lowest. $k$ is configurable to allow for varying the amount of inference-time computation MAV performs. If there are fewer legal moves than $k$, the model simply outputs all of them (this can also be achieved by setting $k$ to ``$all"$). In the case where the starting state is terminal, the %top_k command outputs the game outcome instead (e.g., [%top_1 invalid : "1-0"] when the first player has won).
The action value for a move corresponds to the predicted win probability for the player in the current state if this move is taken. For chess, state-action values are represented similarly to Ruoss et al. (2024a): Stockfish centipawn evaluations are mapped to win probabilities using the formula
@@
obtained from https://lichess.org/page/accuracy. Win probabilities are then mapped to discrete non-overlapping buckets, making the value prediction problem a classification task rather than a regression task. This has proven to be beneficial in other works (Farebrother et al., 2024). We use 64 buckets, each represented by a different special token (e.g., <ctrl28> for bucket 28). Figure 2 shows two win probability distributions predicted by MAV. Connect Four and Hex use game engines Fhourstones (Tromp, ) and neurobenzene (Gao and Pawlewicz, ), respectively, to obtain state-action values, which are similarly mapped to the same 64 bucket tokens.
When producing training data we vary $k$ randomly, making sure to include examples with $k$ greater than the number of legal moves. We randomize the order of the moves to: ($i$) encourage the model to treat the moves independently, and ($ii$) help to prevent hallucinations. With a fixed order, e.g., lexicographical from $a1a2$ to $h8h7$, then ($i$) if the model skipped a legal move $a1a4$ and emitted $a1a5$, it couldn’t go back and fix its mistake, and ($ii$) the model might learn to heavily lean on evaluations of previous moves when considering a move like $h8h7$, which is late in the order. Using randomised ordering of moves in the training data, the model is steered towards an approximate permutation symmetry (Bronstein et al., 2021).
Figure 2: For a safe positional move, such as Re1, MAV assigns little uncertainty and gives a win probability of 40-45% since white is a pawn down. In contrast, the bishop sacrifice Bxh7+ (best move according to Stockfish) is a riskier attacking move and thus has substantial probability mass around low, equal, and high win probabilities.
@@
During inference, for each move the model produces a distribution over the 64 tokens that correspond to action value buckets. We experiment with two methods to turn this distribution into a final score for the move: max scoring, which corresponds to greedy decoding, uses the mode of the model’s distribution, while mean scoring uses the expectation over buckets.
Mean scoring (also used in Ruoss et al. (2024a)) utilizes more of the model’s information, and is better able to distinguish actions with equal modes. In an almost tied position, the most likely outcome for many moves may be a draw, but still some moves may have a higher winning chance than others. Mean scoring will be able to pick the best move in this case, while max scoring can not. This is further illustrated in Figure 2.
@@
In positions where the model has a large advantage, it may have many moves available with ca. 100% win probability that don’t directly lead to finishing the game by checkmate. This may cause the model to play aimlessly in such positions. This problem was encountered in Ruoss et al. (2024a), which they resolved by using Stockfish to break ties when all top 5 moves lie above a win probability of 99%. To eliminate reliance on a game engine at inference time, we introduce the %best_action command, which instructs the model to emit the game engine’s chosen action in the position. In this way we teach the model the tie-breaking procedure described above.
@@
To summarize, MAV makes four key improvements over the existing state-of-the-art Transformer-based chess engines (Ruoss et al., 2024a; Monroe and Leela Chess Zero Team, 2024; Czech et al., 2024; Farebrother et al., 2024) First, it performs world modeling, policy and action-value computation together in one model. Second, it is trained to output the best action at the end of action-value modeling, enabling the model to reliably finish games where it has a decisive advantage. These two improvements enable MAV to play complete games without relying on an external game engine for legal moves or for finishing the game. Third, all of the above steps can be done in a single model call without having to evaluate every action separately—an important feature that guided the design of the MAV format to reduce the cost and infrastructural complexity of inference. Fourth, the amount of compute performed by the model at inference time can be varied dynamically, enabling us to achieve higher quality at the cost of higher latency when performing both internal and external search, where the performance increases as we scale the planning.
@@
We curate a dataset of diverse, relevant positions in four games: Chess, Chess960, Connect Four, and Hex. The statistics and sources for each of these datasets are shown in Table 4 in Appendix B. Each position is used to produce a single training example, randomly varying ($i$) the $k$ action values in %top_k, ($ii$) the presence of the initial or final state tracking commands, ($iii$) the use and order of %state or %FEN representations in chess. Further details on the datasets are provided in Appendix B.
@@
We train two randomly initialized decoder-only Transformer models using the Gemini architecture (Gemini Team et al., 2024), called MAV and MAV–small, [^1] As of May 21, 2025, it is possible to play chess against MAV–small at https://goo.gle/ChessChamp. with 2.7 billion and 1 billion parameters respectively [^2] Note that in practice MAV and MAV–small don’t utilize many natural-language specific parameters.. They were trained on 1.9 and 1.2 epochs of the dataset. Except for the max scoring result for MAV–small in Table 1, all other experiments and results use the 2.7 billion MAV. In order to efficiently use our models’ parameters, the input part of each training example is masked out during loss computation. This means our models do not waste capacity on learning to generate game positions.
@@ 3 External Search
External search employs MAV to generate planning steps, and applies a search algorithm to direct and optimise over sequences of planning steps. In this paper, we evaluate planning on top of the previously discussed MAV, using MCTS as the search procedure.
External search is based on an AlphaZero-style MCTS (Silver et al., 2017). There are two key ingredients: a prior function which returns a probability distribution over actions at each state, and a value function returning a numerical value indicating value of a state or state-action pair, both extracted from the MAV output. External search is an adaptation of MCTS; in its most basic form, as in AlphaZero, MCTS relies on an explicit world model, querying a game engine (e.g., a chess implementation in OpenSpiel (Lanctot et al., 2019)) to provide legal actions, state transitions and terminal states. Inspired by MuZero (Schrittwieser et al., 2020), in a subsection below we describe how we remove this dependency on the game engine and instead use only MAV to track states, transitions, and provide legal actions and game outcome during planning.
@@
The external search algorithm guided by a learned world model is summarized in Algorithm 1, with subroutines contained in Appendix A. A search is started at an initial game state $s_{0}$. First, the legal actions $\boldsymbol{a}_{0}^{L}$ and associated state-action values $\boldsymbol{Q}^{(i)}(s_{0},\boldsymbol{a}_{0}^{L})$ from active player’s $i$ perspective are obtained from MAV. The prior function is an $\varepsilon$-greedy policy, derived entirely from state-action values comprised of a greedy softmax over the top $k$ values, mixed with a uniform distribution over all actions to encourage exploration. Specifically, let $\boldsymbol{a}^{g,k}\subseteq\boldsymbol{a}^{L}$ be the subset of legal actions $\boldsymbol{a}^{L}$ whose values are among the top $k$-ranked values, or equal to $\boldsymbol{a}^{L}$ if $|\boldsymbol{a}^{L}|\leq k$. Define the probability of action $a$ in state $s$ under the greedy policy to be
$\pi^{g,k}(s,a;\boldsymbol{Q}^{(i)})=\left\{\begin{array}[]{ll}\frac{\exp{\frac{1}{\tau}Q^{(i)}(s,a)}}{\sum_{a^{\prime}\in\boldsymbol{a}^{g,k}}\exp{\frac{1}{\tau}Q^{(i)}(s,a^{\prime})}}&\mbox{if $a\in\boldsymbol{a}^{g,k}$};\\ 0&\mbox{otherwise},\end{array}\right.$
@@
The probability of the uniform policy is $\pi^{u}(s,a;\boldsymbol{a}^{L})=\frac{1}{|\boldsymbol{a}^{L}|}$. The prior probability of taking action $a$ in state $s$ is
$\begin{split}P(s,a;\boldsymbol{Q}^{(i)},\boldsymbol{a}^{L},k)&=(1-\varepsilon)\pi^{g,k}(s,a;\boldsymbol{Q}^{(i)})\\ &\quad+\varepsilon\pi^{u}(s,a;\boldsymbol{a}^{L}).\end{split}$ (1)
The root node $\mathcal{N}_{0}$ is then initialized with the active player, string description of a state corresponding to the node, and legal actions (respectively: $\mathcal{N}_{0}(i)$, $\mathcal{N}_{0}(s)$, $\mathcal{N}_{0}(\boldsymbol{a}^{L})$). The node is then expanded so that a child node is added for each action and a prior probability attached to the action leading to the child. Then simulations are run: each simulation starts at the root node, actions and children nodes are selected according to PUCT until a leaf node is expanding and evaluated, and values are backpropagated through the nodes that the simulation visited. After $M$ such simulations, a final move $a^{*}$ is selected.